suppose you make an annual contribution of $5000 to your savings account at the end of each year for five years if your savings account earn 6% interest annualy how much can be withdrawn at the end of five years?
1
Expert's answer
2020-04-16T09:19:34-0400

"5000\times1.06 = 5300" , 1-year


"(5300+5000)\times1.06=10918", 2-year


"(10918+5000)\times1.06 = 16873,08", 3-year

"(16873.08+5000)\times1.06 = 23185,46", 4-year


23185.4648+5000=28185,46


28185.46


no interest is charged in the last year, deducted from the account